# Meta-DAN: towards an efficient prediction strategy for page-level handwritten text recognition

### v1

@@ VIII-B Datasets
Figure VIII-B provides document image examples from the 10 evaluated datasets. As one can note, these datasets cover many page layouts, visual aspects, writing styles, periods and languages.
@@
VIII-C WER results for the comparison of prediction strategies.
Table VIII-C provides the WER scores associated to experiments related to Section VI-E, i.e., when relying on monolingual evaluation and the FCN-M encoder. Interestingly, the WER results do not follow exactly the same trend as that of the CER, by exhibiting a slight advantage for the Faster DAN with reaches an on-average WER of 13.73%, compared to 14.22% for the Meta-DAN. There is no clear explanation for this observation. However, it still remains lower than the DAN WER which reaches 15.15%.

### v2

@@ Appendix A Datasets
Figure A provides document image examples from the 10 evaluated datasets. As one can note, these datasets cover many page layouts, visual aspects, writing styles, periods, and languages.
@@
Line-level splits
Table A details the splits for training, validation, and test used for the different datasets regarding the experiments at line level.
Appendix B Scaling up the DAN encoder
We noticed that the DAN lacked some feature extraction capability when evaluated on challenging datasets such as BRESSAY, which presents some crossed-out text with annotations. We opt for scaling up the DAN encoder only in order to increase the performance while preventing a significant prediction time increase. Indeed, the encoding part of the model is only used once at inference (it is excluded from the iterative decoding stage); so the computations involved are not such significant compared to decoding. The proposed encoder is referred to as FCN-M, and the original one as FCN-S.
Composition of the two versions of the encoder. ConvBlock(X) and DSCBlock(X) indicate the number of feature maps X returned by each convolutional layer. FCN-S [15] FCN-M (ours) Output size layers ConvBlock(16) ConvBlock(32) $H\times W$ ConvBlock(32) ConvBlock(64) $\frac{H}{2}\times\frac{W}{2}$ ConvBlock(64) ConvBlock(128) $\frac{H}{4}\times\frac{W}{4}$ ConvBlock(128) ConvBlock(256) $\frac{H}{8}\times\frac{W}{8}$ ConvBlock(128) ConvBlock(512) $\frac{H}{16}\times\frac{W}{8}$ ConvBlock(128) ConvBlock(512) $\frac{H}{32}\times\frac{W}{8}$ DSCBlock(128) DSCBlock(512) $\frac{H}{32}\times\frac{W}{8}$ DSCBlock(128) DSCBlock(512) $\frac{H}{32}\times\frac{W}{8}$ DSCBlock(128) DSCBlock(512) $\frac{H}{32}\times\frac{W}{8}$ DSCBlock(256) DSCBlock(256) $\frac{H}{32}\times\frac{W}{8}$ # params 1.7 M 17.6 M
The proposed modification mainly consists in doubling the number of feature maps per block while preserving the last one, which conditions the size of the decoder, as detailed in Table B. We use the notation of the original paper [15], where ConvBlock is a Convolution Block made up of three convolutions, instance normalization, and dropout. DSCBlock is equivalent except for the convolutions that are replaced by Depthwise Separable Convolutions [9].
We evaluate these modifications on READ-2016, CASIA-2, BRESSAY, and IAM datasets at line level, and at page level with the DAN. Results are shown in Table B. As one can note, the modified encoder improves the recognition results (or is equivalent) on all the evaluated datasets, whether it is at line or page level. The prediction time is only slightly impacted; we assume this is mainly due to the complexity of learning the reading order. For CASIA-2 and IAM, the prediction time is lower, which can be explained by a reduction of the loop effect (predicting several times the same text part). But for READ-2016 and BRESSAY, the prediction time is higher, which can be the result of the additional computations due to the bigger encoder, and to better recognition (missed text parts are now well recognized, adding extra decoding iterations). The most negative impact is on the READ-2016 dataset with a 6% increase in the prediction time. These results also enhance what has been found in [14]: the page-level DAN outperforms its line-level counterpart while they are based on the same encoder.
Appendix C WER results comparing the prediction strategies.
Table C provides the WER scores associated with the experiments related to Section VI-E, i.e., when relying on the FCN-M encoder.
@@ Appendix B Scaling up the DAN encoder
Interestingly, the WER results do not follow exactly the same trend as that of the CER, by exhibiting a slight advantage for the Faster DAN, which reaches an on-average WER of 13.73%, compared to 14.22% for the Meta-DAN. There is no clear explanation for this observation. However, it still remains lower than the DAN WER, which reaches 15.15%. This would suggest that the Faster-DAN is more reliable when its initial predictions are correct, leading to perfectly recognized words (lower WER), but that bad starts cause cascading errors, leading to higher CER, while the proposed Meta-DAN is more constant in character recognition performance.